A triangle has a perimeter of 28 meters . Find the length of each side of the trianle if the two shorter lenghts multiplied together equal the longer side multiplied by 5
Answer by Nate(3500)

I am guessing that the short lenghts are congruent. (Insuffice)
one short length = s
other short length = t
long side = l
st = 5l which can look like: st/5 = l
s + t + l = 28
s + t + st/5 = 28 remember ~> s = t
s + s + s^2/5 = 28
s^2/5 + 2s - 28 = 0
s^2 + 10s - 140 = 0 You can finish the rest.